# Pinning control of chimera states in systems with higher-order interactions

### v1

@@ V Discussion
In this work we have shown how pinning control can be applied to higher-order system to trigger the emergence of chimera states and how higher-order interactions are a key feature for the chimera state to develop and persist. It was already known from previous works that higher-order interactions enhanced chimera states, however, the set of parameters, initial conditions and couplings allowing for such behavior remained limited. Thanks to our pinning schemes, that we called additive pinning and parametric pinning, we were able to overcome such limitation and observe chimera patterns for a wide range of settings. Moreover, and this is the most remarkable result, the higher-order framework makes possible to control the presence of chimeras by only acting on a small fraction of the nodes, at striking contrast with the network case, where half of the nodes are to be controlled to achieve this objective. Lastly, our heuristic interpretation of the results goes beyond this work and provides a possible explanation of other previous results regarding synchronization patterns observed in higher-order systems, in particular the claim made by Zhang et al. that higher-order interactions shrink the attraction basin of the synchronized state Zhang et al. (2023).
@@
Our results clearly show that it is easier and more efficient to trigger the emergence of chimera states when higher-order interactions are present. A further study could be to determine how much energy is needed to control the chimera state in comparison with the pairwise setting, by relying on energy aware controllability measures Lindmark and Altafini (2018); Baggio et al. (2022). Another efficient control strategy could be to apply an intermittent pinning, analogously to the occasional coupling setting developed in the framework of amplitude death Ghosh et al. (2022), where techniques from piecewise-smooth systems could be used Coraggio et al. (2021). Another interesting direction would be to apply pinning control to directed higher-order structures, such as directed Gallo et al. (1993) and $m$-directed hypergraphs Gallo et al. (2022), which have been proven to greatly affect nonlinear dynamics in the context of synchronization Della Rossa et al. (2023) and Turing pattern formation Dorchain et al. (2024), respectively. Pinning approaches in systems with directed higher-order interactions have been developed in some pioneering works Della Rossa et al. (2023); Shi et al. (2023); Li et al. (2024); Rizzello and De Lellis (2024), but not yet in the context of chimera states. Lastly, the recent implementation of higher-order interactions in electric circuits Vera-Ávila et al. (2024) opens the way to further applications in this direction.
@@
In conclusion, this work is one of the first in which tools from control theory are applied to systems with higher-order interactions and it shows the numerous possibilities offered by this novel framework. We believe that there is plenty of exciting research to be done in this direction and that the ground we built with this work is the basis for further studies shedding more light on the interplay between dynamics with higher-order interactions and its control.
Acknowledgements
@@
R.M. and H.N. acknowledge JSPS, Japan KAKENHI JP22K11919, JP22H00516, and JST, Japan CREST JP-MJCR1913 for financial support. The contribution of L.V.G. and M.F. to this work is framed into the activities of the project ”CoCoS: Control of Complex Systems”, funded by the University of Catania, under the PIA.CE.RI. 2024-26 initiative. We are grateful to Timoteo Carletti for useful discussions, including pointing us to Ref. Shanahan (2010). R.M. is particularly grateful to Giacomo Baggio and Francesco Lo Iudice for the discussions on pinning control and to Iván León and Yuzuru Kato for the discussions on the phase reduction.
@@ Appendix A Numerical results for different coupling schemes
The following configurations lead $40\%-60\%$ of times to chimeras, depending on the parameters $\omega_{i_{p}}$:
@@ Appendix B Scaling with respect to the number of nodes for the parametric pinning control approach
In Fig. B1, we show the results obtained with such control scheme on $d$-hyperrings with $d=3,4,5,6$, i.e., $4$-, $5$-, $6$- and $7$-body interactions, where we have fixed the number of hyperedges. Indeed, we can observe that we obtain a chimera state by inducing a large region of incoherence (more than half of the nodes) with a control that involves only a small fraction of the nodes. Moreover, the pinned nodes $N_{p}$ are kept constant for every structure, meaning that $N_{p}$ does not scale with the number of nodes, but rather with the number of hyperedges, which allows to control large structures with only a handful of nodes. The parameters $\omega_{i_{p}}$ are the same for all the simulations. The pinned nodes are $\simeq 8.8\%$ of the total nodes for the $3$-hyperring (panels a) and b)), $\simeq 6.6\%$ of the total nodes for the $4$-hyperring (panels c) and d)), $\simeq 5.3\%$ of the total nodes in the $5$-hyperring (panels e) and f)) and $\simeq 4.4\%$ of the total nodes in the $6$-hyperring (panels g) and h)). In the latter case, the pinning scheme consisting in controlling one every $2$ junction nodes does not yield a chimera state, not even with stronger coupling, as shown in Fig. B1g-h), where, moreover, we see that the region of incoherence enlarges. This does not change if we increase the number of pinned nodes, but only if we reduce the gap between them.
@@
Figure B1: Scaling of the pinned subset with respect to the higher-order structure size with the parametric pinning approach. On the upper panels we depict the snapshots of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units, while the lower panels show the hyperedge-based local order parameters. Panels a) and b) report the simulations performed on a $3$-hyperring (i.e., $4$-body interactions) of $204$ nodes, panels c) and d) on a $4$-hyperring (i.e., $5$-body interactions) of $272$ nodes, panels e) and f) on a $5$-hyperring (i.e., $6$-body interactions) of $340$ nodes and panels g) and h) on a $6$-hyperring (i.e., $7$-body interactions) of $408$ nodes. The number of nodes is chosen so that each hyperring is made of $68$ hyperedges. Pinning control is applied to $N_{p}=18$ nodes as in Fig. 7, i.e., every two junction nodes, to all the structures and for $t_{p}=100$ time units. This means that the pinned nodes are $\simeq 8.8\%$ of the total nodes in the $3$-hyperring, $\simeq 6.6\%$ of the total nodes in the $4$-hyperring, $\simeq 5.3\%$ of the total nodes in the $5$-hyperring and $\simeq 4.4\%$ of the total nodes in the $6$-hyperring. The model parameters are $\alpha=1$ and $\omega=1$, the coupling strength is $\varepsilon=0.01$, except for the $6$-hyperring where $\varepsilon=0.1$, and the parameters $\omega_{i_{p}}$ are drawn from a uniform distribution in the interval $[1.5,5.5]$ and are the same for all the simulations.
Appendix C Parametric pinning control of the reduced phase model
In this Appendix, we complement the numerical results of Sec. IV of the Main Text on additive pinning control for the phase reduced model by showing analogous results for the parametric pinning approach. In Figs. C1 and C2, we show the results of the parametric pinning procedure on a $3$-hyperring and its corresponding clique-projected network, respectively. The setting is analogous to that of Figs. 5 and 6 of the Main Text. We see that, while the control induces a chimera state in the higher-order phase reduced model, only a slight incoherence emerges in the pairwise setting, which is not a chimera, as can be seen by looking at the clique-based local order parameter.
Figure C1: Parametric pinning induces phase chimera states for the reduced phase model (17) of $204$ nodes. Panel a) depicts the whole time series of the phases $\vartheta_{i}(t)$ with $i=1,...,N$, panel b) the hyperedge-based local order parameter, panel c) a snapshot of the phases $\vartheta_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of the phases $\vartheta_{i}(t)$ with $i=1,...,N$. The frequency is $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\omega_{i_{p}}$ are the same of Figs. 5 and 6.
Figure C2: Parametric pinning on a clique-projected network of $204$ nodes. Chimera states do not emerge in this setting, at contrast with the previous Fig. C1. Nonetheless, there appears to be a region of weak incoherence, but, as shown by the local order parameter, it is not a chimera because the oscillators remain synchronized with their neighbors. Panel a) depicts the whole time series of the phases $\vartheta_{i}(t)$ with $i=1,...,N$, panel b) the clique-based local order parameter, panel c) a snapshot of the phases $\vartheta_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units and panel d) shows a zoom of the time series of the phases $\vartheta_{i}(t)$ with $i=1,...,N$. The model parameters are $\alpha=1$ and $\omega=1$ and the coupling strength is $\varepsilon=0.01$. Pinning control is applied to $N_{p}=40$ consecutive nodes for $t_{p}=100$ time units. The parameters $\omega_{i_{p}}$ are the same of Figs. 5 and 6.

### v2

@@ V Discussion
In this work we have shown how pinning control can be applied to higher-order systems to trigger the emergence of chimera states and how higher-order interactions are a key feature for the chimera state to develop and persist. It was already known from previous works that higher-order interactions enhanced chimera states, however, the set of parameters, initial conditions and couplings allowing for such behavior remained limited. Thanks to our pinning schemes, that we called additive pinning and parametric pinning, we were able to overcome such limitations and observe chimera patterns for a wide range of settings. Moreover, and this is the most remarkable result, the higher-order framework makes possible to control the presence of chimeras by only acting on a small fraction of the nodes, at striking contrast with the network case, where about half of the nodes need to be controlled to achieve this objective. Lastly, our heuristic interpretation of the results goes beyond this work and provides a possible explanation of other previous results regarding synchronization patterns observed in higher-order systems, in particular the claim made by Zhang et al. that higher-order interactions shrink the attraction basin of the synchronized state zhang2024deeper .
@@
Our results clearly show that it is easier and more efficient to trigger the emergence of chimera states when higher-order interactions are present. A further study could be to determine how much energy is needed to control the chimera state in comparison with the pairwise setting, by relying on energy aware controllability measures lindmark2018minimum ; baggio2022energy . Indeed, in zhang2024deeper , as well as in other works such as von2024higher , it is proven that the synchronous state is more robust to perturbations in the presence of higher-order interactions. This means that, in principle, it may be more costly, in terms of energy, to desynchronize coupled oscillators on hypergraphs. A further direction of study could be an estimation from the energy needed to obtain chimera states in our setting. They are easier to obtain, with respect to the case of networks, and with a longer life time, but how costly are they? This question may be very challenging, because the main energy estimation methods developed in control theory consist in computing the energy needed to steer the system onto a desired trajectory from a given initial state baggio2022energy ; in the case of chimera states, we do not have an expression of the desired trajectory, not even numerically, making the estimation problem difficult. Another efficient control strategy could be to apply an intermittent pinning, analogously to the occasional coupling setting developed in the framework of amplitude death ghosh2022occasional , where techniques from piecewise-smooth systems could be used coraggio2021convergence . Another interesting direction would be to apply pinning control to directed higher-order structures, such as directed gallo1993directed and $m$-directed hypergraphs gallo2022synchronization , which have been proven to greatly affect nonlinear dynamics in the context of synchronization della2023emergence and Turing pattern formation dorchain2024impact , respectively. Pinning approaches in systems with directed higher-order interactions have been developed in some pioneering works della2023emergence ; shi2023synchronization ; li2024synchronization ; rizzello2024pinning , but not yet in the context of chimera states. Lastly, the property of higher-order structures to conserve the incoherence could be exploited by engineering the perturbation to attain a desired incoherent state. A possible application could be in the framework of electric circuits, where higher-order interactions have been recently implemented vera2024electronic ; minati2024chaotic .
@@
In conclusion, this work is one of the firsts in which methods from control theory are applied to systems with higher-order interactions and it shows the numerous possibilities offered by this novel framework. We believe that there is plenty of exciting research to be done in this direction and that the ground we built with this work sets the basis for further studies shedding more light on the interplay between dynamics with higher-order interactions and its control.
Acknowledgements.
@@
The work of R.M. is supported by a JSPS postdoctoral fellowship, grant 24KF0211. H.N. acknowledges JSPS KAKENHI 25H01468, 25K03081, and 22H00516 for financial support. The contribution of L.V.G. and M.F. to this work is framed into the activities of the project ”CoCoS: Control of Complex Systems”, funded by the University of Catania, under the PIA.CE.RI. 2024-26 initiative. We are grateful to Timoteo Carletti for useful discussions, including pointing us to Ref. shanahan2010metastable . R.M. is particularly grateful to Martin Moriamé, Giacomo Baggio and Francesco Lo Iudice for the discussions on pinning control and to Iván León and Yuzuru Kato for the discussions on the phase reduction.
Conflict of interest
The authors declare that they have no conflict of interest.
Author contributions
Conceptualization: R.M., L.V.G., M.F.; Methodology: R.M., L.V.G., H.N., M.F.; Software: R.M.; Formal Analysis: R.M., H.N.; Validation: R.M.; Writing-Original Draft: R.M.; Writing-Review and Editing: R.M., L.V.G., H.N., M.F.; Funding Acquisition: H.N., M.F.
@@ Appendix A Numerical results for different coupling schemes
The following configurations lead $40\%-60\%$ of times to chimeras, while the remaining times to weak chimera-like states, depending on the parameters $\omega_{i_{p}}$:
@@ Appendix B Scaling with respect to the number of nodes for the parametric pinning control approach
In Fig. B1, we show the results obtained with such control scheme on $d$-hyperrings with $d=3,4,5,6$, i.e., $4$-, $5$-, $6$- and $7$-body interactions, where we have fixed the number of hyperedges. Indeed, we can observe that we obtain a chimera state by inducing a large region of incoherence (more than half of the nodes) with a control that involves only a small fraction of the nodes. Moreover, the pinned nodes $N_{p}$ are kept constant for every structure, meaning that $N_{p}$ does not scale with the number of nodes, but rather with the number of hyperedges, which allows to control large structures with only a handful of nodes. The parameters $\omega_{i_{p}}$ are the same for all the simulations. The pinned nodes are $\simeq 8.8\%$ of the total nodes for the $3$-hyperring (panels a) and b)), $\simeq 6.6\%$ of the total nodes for the $4$-hyperring (panels c) and d)), $\simeq 5.3\%$ of the total nodes in the $5$-hyperring (panels e) and f)) and $\simeq 4.4\%$ of the total nodes in the $6$-hyperring (panels g) and h)). In the latter case, the pinning scheme consisting in controlling one every $2$ junction nodes does not yield a chimera state, not even a weak chimera-like state, as shown in Fig. B1g-h), where, moreover, we see that the region of incoherence enlarges. This does not change if we increase the number of pinned nodes, nor increase the coupling strength, but only if we reduce the gap between them.
@@
Figure B1: Scaling of the pinned subset with respect to the higher-order structure size with the parametric pinning approach. On the upper panels we depict the snapshots of variables $y_{i}(t)$ with $i=1,...,N$ for $t_{final}=1000$ time units, while the lower panels show the hyperedge-based local order parameters. The results for the variables $x_{i}(t)$ are analogous and, hence, not shown. Panels a) and b) report the simulations performed on a $3$-hyperring (i.e., $4$-body interactions) of $204$ nodes, panels c) and d) on a $4$-hyperring (i.e., $5$-body interactions) of $272$ nodes, panels e) and f) on a $5$-hyperring (i.e., $6$-body interactions) of $340$ nodes and panels g) and h) on a $6$-hyperring (i.e., $7$-body interactions) of $408$ nodes. The number of nodes is chosen so that each hyperring is made of $68$ hyperedges. Pinning control is applied to $N_{p}=18$ nodes as in Fig. 7, i.e., every two junction nodes, to all the structures and for $t_{p}=100$ time units. This means that the pinned nodes are $\simeq 8.8\%$ of the total nodes in the $3$-hyperring, $\simeq 6.6\%$ of the total nodes in the $4$-hyperring, $\simeq 5.3\%$ of the total nodes in the $5$-hyperring and $\simeq 4.4\%$ of the total nodes in the $6$-hyperring. The model parameters are $\alpha=1$ and $\omega=1$, the coupling strength is $\varepsilon=0.01$, except for the $6$-hyperring where $\varepsilon=0.1$, and the parameters $\omega_{i_{p}}$ are drawn from a uniform distribution in the interval $[1.5,5.5]$ and are the same for all the simulations.